OpenStudy (anonymous):

Solve for p: 3p - 2(p-4) = 7p + 6

OpenStudy (math&ing001):

Want me to walk you through ?

OpenStudy (anonymous):

yeah

OpenStudy (math&ing001):

Ok first distribute this part -2(p-4)

OpenStudy (math&ing001):

You know how to distribute?

OpenStudy (anonymous):

not really I just started and I barely get anything

OpenStudy (math&ing001):

Ok here's how a(b+c) = ab + ac Try applying it to our term -2(p-4).

OpenStudy (math&ing001):

Don't worry if you get it wrong I just want you to try.

OpenStudy (anonymous):

-2*p--2*4?

OpenStudy (math&ing001):

Yeah correct ! can you simplify it further, knowing that: --=+

OpenStudy (anonymous):

-2*p+2*4?

OpenStudy (math&ing001):

Yeah and 2*4=8, so we'll have 3p - 2p +8 = 7p + 6. Clear till now ?

OpenStudy (anonymous):

yep

OpenStudy (math&ing001):

Now we're gonna move all terms that contain p to one side, and the rest to the other side. If you subtract 7p from both sides what would you get?

OpenStudy (anonymous):

what do you mean by both sides? what are the two sides?

OpenStudy (math&ing001):

3p - 2p +8 = 7p + 6. 1st side = 2nd side

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ist side=2p 2nd=6p ?

OpenStudy (math&ing001):

Now do you think you can do that, subtract 7p from both sides

OpenStudy (math&ing001):

No you have to keep the equation intact, I'll do this one for you. 3p - 2p +8 = 7p + 6 -7p -7p 3p - 2p - 7p +8 = 6

OpenStudy (anonymous):

oh I think I get it!

OpenStudy (math&ing001):

Ok if you get it, now it's your turn. Subtract 8 from both sides.

OpenStudy (anonymous):

the rest I can do I just didn't know how to start, thanks so much for your help

OpenStudy (math&ing001):

Welcome !